OpenStudy (anonymous):

find the limit of the sequence a_n = (1/(e^(4n)+n^2)^(1/n)?

OpenStudy (anonymous):

sequence goes to infinity

OpenStudy (anonymous):

\[\lim_{n\to \infty}\frac{1}{\left(e^{4n}+n^2\right)^{\frac{1}{n}}}\]?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i think the best bet is to rewrite as \[\left(e^{4n}+n^2\right)^{-\frac{1}{n}}\] then take the log to get \[-\frac{1}{n}\ln(e^{4n}+n^2)\] and use l'hopital

OpenStudy (anonymous):

is -ln(e^(4n)+n^2) the same as ln(e^(4n)+n^2)^-1?